find the zeros of the quadratic function using the square root method h(x)=(x+6)^2-36
Answer by lwsshak3(11628)

find the zeros of the quadratic function using the square root method h(x)=(x+6)^2-36
...
(x+6)^2-36=0
(x+6)^2=36
take the square root of both sides
x+6=±6
x=-6±6
x=0
or x=-12
..
and alternate method is to use different of squares
(x+6)^2-36=0
(x+6+6)(x+6-6)=0
(x+12)(x)=0
x=-12 or 0